OpenStudy (anonymous):

A kid saves c cents per day in a piggy bank. When he opened the bank, he counted a total of 325 dollars. Find the number of weeks he saved for the total amount. (Use ratio/proportion)

OpenStudy (darthsid):

People in the Mathematics group can help you much better than me, for such questions. I will, however, give it a shot. Total number of cents : 32500 Cents saved per day: c Total number of days : \[\frac{32500}{c}\] Days in a week: 7 Total number of weeks: \[\frac{\frac{32500}{c}}{7}\] Simplify that to obtain your answer!